OpenStudy (anonymous):

A spherical balloon is being inflated so that its volume is increasing at the rate of 5m^3/min. At what rate is the Diameter increasing when the diameter is 12m?

OpenStudy (anonymous):

i know this: dv/dt=5m^3/min d=12 and radius= 6m Formula of sphere: V=4/3pi(r)^3

OpenStudy (anonymous):

then I got this: dr/dt= 5/144pi

OpenStudy (anonymous):

I'm not sure how to get the rate of the diameter... please help...

OpenStudy (mathmath333):

\(d\color{red}{r}/dt\) u have found. but u need to find \(d\color{red}{d}/dt\) as the rate of diameter is asked not radius

OpenStudy (anonymous):

i know, I'm sure how to get dd/dt

OpenStudy (mathmath333):

\(\large\color{black}{\begin{align} V&=\dfrac{4}{3}\pi(r)^3\hspace{.33em}\\~\\ V&=\dfrac{4}{3}\pi(\dfrac{d}{2})^3\hspace{.33em}\\~\\ V&=\dfrac{4}{3}\pi\dfrac{d^3}{8}\hspace{.33em}\\~\\ V&=\dfrac{1}{6}\pi\cdot d^3\hspace{.33em}\\~\\ \dfrac{dV}{dt}&=\dfrac{1}{6}\pi\cdot \dfrac{d'd^3}{dt}\hspace{.33em}\\~\\ 5&=\dfrac{1}{6}\pi\cdot \dfrac{dd^3 }{d'd'}\dfrac{d'd'}{dt}\hspace{.33em}\\~\\ 5&=\dfrac{1}{6}\pi\cdot3d^2\dfrac{d'd'}{dt}\hspace{.33em}\\~\\ 5&=\dfrac{1}{2}\pi\cdot12^2\dfrac{d'd'}{dt}\hspace{.33em}\\~\\ 5&=72\pi\cdot\dfrac{d'd'}{dt}\hspace{.33em}\\~\\ \dfrac{d'd'}{dt} =\dfrac{5}{72\pi}\quad m/min\hspace{.33em}\\~\\ \end{align} }\)

OpenStudy (mathmath333):

i have just used \('\quad '\) to seperate the two \(d\)

OpenStudy (anonymous):

okay! i got it. Thank you!

OpenStudy (mathmath333):

u got \(dr/dt= 5/144\pi\) by just multiplying it by \(2\) u could have reached \(2dr/dt= 5/72\pi\\ dd/dt= 5/72\pi\\ \)

OpenStudy (anonymous):

what do you mean by multiplying?

OpenStudy (anonymous):

@mathmath333